These 3 problems are from a worksheet, not a textbook, and I so desperately need them solved within the next hour, if possible. I tried and cannot work them out. Thanks!
1.Solve the following exponential equation. Your solution should be accurate to 2 decimal places.
41/(1-1e^7x) = 56




2. Solve the following exponential equation. Your solution should be accurate to 2 decimal places.
5e^-5 + 5 = 7



3. Solve the following logarithmic equation. Your solution should be accurate to 2 decimal places.
ln(x-4) + ln(7) = 4
Thank you very much for your help with this tonight.


Answer by scott8148(5896)

1) 41/56=1-1e^7x

41/56-1=-1e^7x

15/56=e^7x

ln(15/56)=7x

(ln(15/56))/7=x


2) no unknown


3) (x-4)7=e^4

x-4=(e^4)/7

x=((e^4)/7)+4